OpenStudy (anonymous):

Using the following Exam Scores and Frequencies, construct a cumulative frequency distribution. Include the relative frequency and midpoint as well as your cumulative frequency. Exam Scores Frequency 41-50 2 51-60 2 61-70 4 71-80 8 81-90 9 91-100 5

OpenStudy (anonymous):

Will medal & fan!!!

OpenStudy (nurali):

41-50 2 2 51-60 2 2+2=4 61-70 4 4+4=8 71-80 8 8+8=16 81-90 9 16+9=25 91-100 5 25+5=30

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

@Nurali ?

OpenStudy (nurali):

cumulative frequency. 2 2+2=4 4+4=8 8+8=16 16+9=25 25+5=30